OpenStudy (anonymous):

Toby deposited $2,103.24 in a savings account that earns 5.1% simple interest. What will Toby’s account balance be in 10 months? $2,210.51 $2,103.24 $2,192.63 $3,175.89

OpenStudy (ghazi):

(2103.24 *5.1* 0.833)/100= 89.38.....therefore balance= 2103.24+89.38= $2192.63